Find the value of y for the given value of x by first substituting the value of x into the equation and solving for y.


Found 2 solutions by rfer, dyakobovitch:
Answer by rfer(16322)

4x-3xy=16
4(-2)-3(-2)(y)=16
-8+6y=16
subtract +8 from both sides to isolate the 6y
-8+6y+8=16+8
6y=24
divide both sides by 6
6y/6=24/6
y=4
--------------------
When doing this problem watch the signs very closely.

Answer by dyakobovitch(40)

Here we are dealing with a classic case of pure substitution. Simply take your value for x and substitute it in to find y.

4x-3xy=16; x=-2; 4(-2) -3(-2)y=16
So, -8 +6y =16.
6y=24.
y=4

Feel free to leave comments with any questions or concerns. -- David